OpenStudy (word2):

F = Gm^2/r^2 . solve for G

OpenStudy (word2):

@mathmate

OpenStudy (mathmate):

What have you done so far?

OpenStudy (word2):

so i know that you have take the MULTIPLY the r^2 to both sides so now i'm left with Fr^2 =Gm^2 but i don't know what else to do

OpenStudy (word2):

so maybe its G= Fr^2/m^2 ?

OpenStudy (word2):

or is it Fm^2/ r^2

OpenStudy (mathmate):

What you could do with multiply could be done with divide as well. Fr^2=Gm^2 So there you are, dividing both sides by m^2 gives Fr^2/m^2=G or G=Fr^2/m^2

OpenStudy (word2):

i'm so confused. so from F = Gm^2/r^2 . why wouldn't you multiply m^25 to both sides to get Fm^ 2 = G/r^2 , and then do the same with r^2 GIVING YOU Fm^2/r^2

OpenStudy (photon336):

@word2 When we say solve for G we mean that we have to get G by itself \[F = \frac{ G~m^{2} }{ r^{2} }\] the first thing we need to do is multiply both sides by r^2 \[F*(r^{2}) = \frac{ G~m^{2} }{ \cancel\r^{2} }*(\cancel\r^{2})\] \[F*(r^{2}) = G~m^{2}\] What do you think you should do now?

OpenStudy (word2):

take out the G, so divide it to the other side